Using substitution method:4x-5y=39,-2x+7y=-51
Answer by ankor@dixie-net.com(22740)

Using substitution method:
4x-5y=39,
-2x+7y=-51
:
Use the 2nd equation for substitution
-2x = -7y - 51
divide both sides by -2
x = 3.5y + 25.5
:
Substitute (3.5y+25.5) for x in the 1st equation
4(3.5y+25.5) - 5y = 39
14y + 102 - 5y = 39
14y - 5y = 39 - 102
9y = -63
y = -63/9
y = -7
:
Find x, replace y with -7
x = 3.5(-7) + 25.5
x = -24.5 + 25.5
x = 1
:
:
Check solutions in the 1st equation: 4x-5y=39,
4(1) - 5(-7) = 39
4 + 35 = 39